OpenStudy (anonymous):

Please help me with this question: a force of 50N accelerates a 5.0kg block at 6.0m/s^2 along a horizontal surface. What is the frictional force acting on the block?

OpenStudy (anonymous):

Alright so Force F is given by F = mass * acceleration If a mass of 5 kg was accelerated across a frictionless surface by a force of fifty newtons then acceleration would have been given by F/mass = acceleration 50/5 = 10 m s^-2 But as this is not the actual acceleration a frictional force must be at work. Calculating this force is as easy as calculating the acting force (50 N) and then calculating what the what force it would take to accelerate the specified mass to the specified speed which is again the equation F = mass * acceleration F = 5 kg * 6 m s^2 F = 30 N The frictional force then equals the acting force minus the theoretical frictionless force: 50 N - 30 N = 20 N Which is to say the frictional force acting on the block equals 20 N

OpenStudy (abhisar):

Good work @Garadon_Shagan !

OpenStudy (anonymous):

Thanks @Abhisar ! :D

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

@ernesto22 \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)